OpenStudy (anonymous):

Which trinomial is the product of the binomials (4n + 5m)(3n – m)? A. 7n^² + 4mn – 5m^² B. 12n^² + 11mn – 5m^² C. 12n^² + 11mn + 5m^² D. 12n^² + 19mn – 5m^²

OpenStudy (anonymous):

do you know how to do this? you need to do four multiplications

OpenStudy (anonymous):

what is \(4n\times 3n\)?

OpenStudy (anonymous):

12n

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

\(4\times 3=12\) but \(n\times n=n^2\) so it is \(4n\times 3n=12n^2\) for the first term

OpenStudy (anonymous):

that eliminates answer A

OpenStudy (anonymous):

how about \(5m\times (-m)\)?

OpenStudy (anonymous):

4m

OpenStudy (anonymous):

no, you are not adding, you are multiplying

OpenStudy (anonymous):

Is it 5m^-2

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-5m^2

OpenStudy (anonymous):

oops, that eliminates C

OpenStudy (anonymous):

I was kinda close, just had my negative in the wrong spot

OpenStudy (anonymous):

right two more \(4n\times (-m)+5n\times 3m\)

OpenStudy (anonymous):

It b.

OpenStudy (anonymous):

So it would be B?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

yw